### v1

@@ Abstract
We present a novel implementation of a soft sphere, discrete elements code to simulate the dynamics of self-gravitating granular materials. The code is used to study the outcome of sub-sonic collisions between self-gravitating rubble piles with masses ranging from $\sim 6\times 10^{21}$ to $\sim 6\times 10^{22}$ g. These masses are representative of asteroids and planetesimals in the $\sim 100$ km range. We simulate rubble piles composed of a range of particle sizes and analyze the collisions outcome focusing on the properties of the largest surviving fragment. We successfully test and validate the code against previous results. The results of our study show that rubble piles formed by collision of two parent rubble piles do not maintain the same particle size distribution as their parents. Rubble piles formed in low velocity collisions are characterized by a larger fraction of large particles, while the largest fragments of high-velocity collisions show a significant decrease in their mean particle size. We ascribe this effect to the fact that large particles transmit most of the forces during the collisions. In addition, we find that the mass of the largest post-collision fragment does depend on the rotation of the colliding rubble piles. This effect is especially noticeable when the pre-collision spin axes are parallel with each other and perpendicular to the relative velocity. This finding can be particularly relevant for meter to kilometer sized bodies embedded in protostellar accretion disks, where viscous stresses can efficiently align the target and projectile spin axes.
@@ 1 Introduction
While most of the baryonic matter in the Universe is in gaseous and plasma phase within stars, galaxies, and the intergalactic medium (Copi et al., 1995; Péroux & Howk, 2020), solids do play many crucial roles in the structure and evolution of the Cosmos (Draine, 2003). Solid phase materials are commonly found in the interstellar medium (ISM) as dust grains (Mathis, 1990; Weingartner & Draine, 2001; Calzetti et al., 2000) and in planetary systems as rocky planets and minor bodies, such as asteroids, comets, etc (Lissauer, 1993; Chapman et al., 1978; Walsh, 2018; Spinrad, 1987). Dust grains are known to have an effect on the thermodynamics (Spitzer, 1978) of the ISM and to be receptacles for interstellar chemistry (Duley & Williams, 1981; Hasegawa & Herbst, 1993). Collisions among grains lead to the formation of larger objects that can grow to become asteroids, planetesimals, and eventually terrestrial planets (Goldreich & Ward, 1973; Johansen et al., 2007; Blum & Wurm, 2008; Wyatt, 2008). Some of the solids in space are whole, single pieces, while others — like dust conglomerates or asteroids and planetesimals — are clusters made by loosely bound individual units or particles (Walsh, 2018). The larger clusters are held together mainly by gravity, while the small ones are dominated by contact forces.
Symbol           Description Units
$R_{i}$ Radius of the $i$-th particle cm
$m_{i}$ Mass of the $i$-th particle cm
@@
$\vec{v}_{i}$ Velocity vector if the $i$-th particle cm/s
$\vec{\omega}_{i}$ Angular velocity vector if the $i$-th particle rad/s
@@
$\vec{F}_{\rm{elastic}}$ The normal force between two particles due to the elatic reaction between the two dyne
$\vec{F}_{\rm{i,friction}}$ The kinetic friction force on particle $i$ dyne
$\vec{F}_{\rm{i,rolling\,friction}}$ The rolling friction force on particle $i$ dyne
$\vec{T}_{\rm{ij}}$ The torque on particle $i$ due to its interaction with particle $j$ dyne cm
$k$ Spring constant dyne/cm
$\mu_{k}$ Coefficient of kinetic friction
$\mu_{r}$ Coefficient of rolling friction
$\delta{t}$ time step of the code s
@@
In general, based on observational evidence and previous studies (Wetherill, 1967; Dohnanyi, 1969; Chapman et al., 1978; Farinella et al., 1982; Davis et al., 1985; Michel et al., 2001; Benavidez et al., 2012; Leinhardt & Stewart, 2012; Walsh, 2018), the outcome of collisions between clusters is not always constructive. Here, by constructive we mean that the largest debris of the collision is more massive than the target cluster [^1] Here and in the following we will call the projectile the cluster with the lower mass, while we will call the target the more massive one. Since our calculations are carried out in the center of mass frame, in which the net momentum is zero, the projectile also has the largest velocity. In case the two rubble piles have the same mass, one of the two will arbitrarily be the projectile and the other the target, if such distinction is needed.. While it is expected that higher velocity collisions have a higher likelihood of being destructive, the collision outcome is not simply predictable by the energy balance of the collision. Previous work (Farinella et al., 1982; Love & Ahrens, 1996; Asphaug et al., 1998; Durda et al., 2004; Richardson et al., 2009) has indeed shown that even collisions with larger kinetic energy than binding energy (in absolute value) can be constructive, resulting in a massive, bound cluster surrounded by fast debris that carry most of the excess kinetic energy. Additional energy is dissipated in non-elastic collisions that increase the thermal energy of the particles. Leinhardt & Stewart (2012), for example, found that even collisions in which the center of mass (CoM) kinetic energy is three times larger than the gravitational potential binding energy of the clusters can be constructive. The high resilience of granular objects (or clusters) to destructive collisions can be ascribed to two main properties that differentiate them from gaseous bodies. First, clusters of hard particles obey the dynamics of granular materials, in which forces are not due to pressure gradients but to contact stresses, and form filamentary strains that propagate the impact in a disproportionate way to a few particles in the cluster, leaving the majority only marginally affected (Sánchez & Scheeres, 2011, 2012, 2014). This creates the ”fast debris” effect in which a few particles acquire high speed from the collision, in greater excess from the escape velocity. These fast particles absorb the incoming kinetic energy, allowing the bound remnant to grow more massive than the initial target cluster. In addition, since the forces generated by the collisions are propagated through the clusters via contact points, the non-elasticity of the particles that compose the cluster allow them to absorb a fraction of the impact energy at every contact point Tanga et al. (2009); Ferrari & Tanga (2020). This allows for some of the impact energy to be converted into internal energy (hotter particles).
@@
In this paper we present a novel code to study the outcome of subsonic collisions between clusters with masses in the asteroid and planetesimal range. We first present the code (Section 2) and compare its results with the outcome of previous studies (Section 3.1). We then use the capabilities of our code to study the effect of collisions on the particle size distribution of the pre- and post-collision clusters, as well as the effect of rotation of the projectile and target clusters (Section 3). We also analyze the dynamic of the collisions to tease out which particle size is mostly involve in propagating forces during the impact. We finally discuss our results and compare with previous studies (Section 4).
@@ 2 The DECCO Code
In this section we present our soft-sphere, discrete element code DECCO (Discrete Element Cosmic COllision) that we used in the calculations presented in section 3. The code uses the soft-sphere technique to compute the binary collision between particle pairs [^2] Collisions between multiple particles are also considered as the cumulative effect of multiple binary collisions involving the same particles., and includes the effects of non-elastic collisions as well as kinetic and rotational friction between particles. At the present stage of development, DECCO only includes gravity as the attractive force and is therefore best suited for collisions between macroscopic bodies, such as rubble-pile asteroids or planetesimals. Since DECCO is based on individual particle-particle interactions, it is applicable for impact velocities that are lower than the speed of sound in the particles. Hydrodynamic codes would be needed for faster collisions (see, e.g., Michel et al. 2001; Leinhardt & Stewart 2012; Michel et al. 2020). DECCO also calculates explicitly all binary interaction, yielding increased accuracy at the cost of $O(n^{2})$ complexity.
@@ 2.1 Free particles
When there is no contact between two particles, i.e., the distance from their centers is larger than the sum of the two particle radii ($r_{ij}>Ri+R_{j}$), the only acting force is gravitational attraction. In this case the force that the $i$-th particle feels because of the presence of the $j$-th particle is:
@@
Figure 2: Example outcome of the post-collision sub-cluster identification algorithm. The figure shows the post-collapse configuration from an $\eta=3$ collision between two identical rubble piles (5000 particles each). The area shown is where the largest debris are found. Particles belonging to dynamically bound sub-clusters with at least 50 particle members are color coded. Dark gray particles are free and unbound. There are six sub-clusters identified in the figure. The two biggest ones (blue and green) are almost identical due to the symmetry of the collision.
@@ 2.2 Radial contact forces
The implementation of DECCO presented here includes the elastic repulsion between two touching particles as the only radial contact force. To take into account the non-elastic behaviour of real materials, DECCO utilizes two different values of Hook’s constant: a larger value $k_{\rm{in}}$ for the incoming trajectory (particles approaching each other) and a smaller value $k_{\rm{out}}$ for the outgoing trajectory. It can be easily shown that this methodology ensures a constant coefficient of restitution $\epsilon=\sqrt{k_{\rm{out}}/k_{\rm{in}}}$.
@@
$\vec{F}_{\rm{elastic}}=k\frac{R_{i}+R_{j}-||\vec{r}_{ij}||}{2}\hat{r}_{ij}$ (3)
where $||\vec{r}_{ij}||$ is the distance between particles $i$ and $j$, $R_{i}$ and $R_{j}$ are the respective radii of particles $i$ and $j$, and the factor $\frac{R_{i}+R_{j}-||\vec{r}_{ij}||}{2}$ measures the compression experienced by particle $i$ from particle $j$. When particles are colliding, this additional surface normal force is added to the gravitational force for the current timestep to simulate the collision ‘bounce’.
@@ 2.3 Restitution
As introduced above, Kinetic energy is removed from particles in collision by a coefficient of restitution directly applied through a change in the elasticity constant, $k$. This method allows for scale-free, constant restitution, which is particularly important since we aim at simulating objects made by multiple particles of different sizes.
@@ 2.4 Friction
$\vec{v}_{i,surface}=\underbrace{\vec{v}_{i}-(\vec{v}_{i}\cdot\vec{r}_{ij})\frac{\vec{r}_{ij}}{||\vec{r}_{ij}||}}_{\text{velocity component tangent to surface}}+\underbrace{\vec{\omega}_{i}\times\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}}_{\text{velocity due to rotation}}$ (4)
@@
$\vec{F}_{i,friction}=-\mu_{k}||\vec{F}_{\rm{elastic}}||\hat{v}_{i,surface,relative}$ (6)
where $\mu_{k}$ is the coefficient of friction.
In addition to sliding friction, the code includes rolling friction, which is calculated as:
$\vec{F}_{i,rollingFriction}=-\mu_{r}||\vec{F}_{\rm{elastic}}||\frac{(\vec{\omega}_{i}-\vec{\omega}_{j})\times\vec{r}_{ij}}{||(\vec{\omega}_{i}-\vec{\omega}_{j})\times\vec{r}_{ij}||}$ (7)
where $\mu_{r}$ is the coefficient of rolling friction. It has to be noted that modeling rolling friction is a complicated task since any deviation from pure spherical shape of the particles would affect their rolling behavior much more than their sliding behavior. To understand the reason, let us consider a small particle rolling on a big one. If the small particle is non spherical, the rolling motion would cause the distance between the two centers of mass to change with time, a kind of motion that requires more energy than the rolling of spherical particles that does not involve any changes in distance.
The total torque on particle $i$ due to the contact with particle $j$ is then:
@@
and the equation for the evolution of the angular velocity is then:
$\frac{2}{5}m_{i}R_{i}^{2}\frac{d\omega_{i}}{dt}=\sum_{i\neq j}\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}\times(\vec{F}_{i,friction}+\vec{F}_{i,rollingFriction})$ (9)
The friction force is equal and opposite for particle $j$, so the necessary information to move both interacting particles is known from a single calculation. By dividing the torque by the respective moment of inertia, the angular acceleration applied due to surface friction is found.

### v2

@@ Abstract
We present a new implementation of a soft-sphere discrete element code to simulate the dynamics of self-gravitating granular materials. The code is used to study the outcome of sub-sonic collisions between self-gravitating rubble piles with masses ranging from $\sim 6\times 10^{21}$ to $\sim 6\times 10^{22}$ g. These masses are representative of asteroids and planetesimals in the $\sim 100$ km range. We simulate rubble piles composed of a range of particle sizes and analyze the collision outcomes focusing on the properties of the largest surviving fragment. We successfully test and validate the code against previous results. The results of our study show that rubble piles formed by collisions of two parent rubble piles do not maintain the same particle size distribution as their parents. Rubble piles formed in low velocity collisions are characterized by a larger fraction of large particles, while the largest fragments of high-velocity collisions show a decrease in their mean particle size. In both cases the effect is small, but could build up to a noticeable difference after multiple collisions. We ascribe this effect to the fact that large particles transmit most of the forces during the collisions. In addition, we find that the mass of the largest post-collision fragment depends on the rotation of the colliding rubble piles. This effect is especially noticeable when the pre-collision spin axes are parallel with each other and perpendicular to the relative velocity. This finding can be particularly relevant for meter to kilometer sized bodies embedded in protostellar accretion disks, where viscous stresses can efficiently align the target and projectile spin axes.
@@ 1 Introduction
While most of the baryonic matter in the Universe is in gaseous and plasma phase within stars, galaxies, and the intergalactic medium (Copi et al., 1995; Péroux & Howk, 2020), solids do play many crucial roles in the structure and evolution of the Cosmos (Draine, 2003). Solid phase materials are commonly found in the interstellar medium (ISM) as dust grains (Mathis, 1990; Weingartner & Draine, 2001; Calzetti et al., 2000) and in planetary systems as rocky planets and minor bodies, such as asteroids, comets, etc (Lissauer, 1993; Chapman et al., 1978; Walsh, 2018; Spinrad, 1987). Dust grains are known to have an effect on the thermodynamics (Spitzer, 1978) of the ISM and to be receptacles for interstellar chemistry (Duley & Williams, 1981; Hasegawa & Herbst, 1993). Collisions between grains lead to the formation of larger objects that can grow to become asteroids, planetesimals, and eventually terrestrial planets (Goldreich & Ward, 1973; Johansen et al., 2007; Blum & Wurm, 2008; Wyatt, 2008). Some of the solids in space are whole, single pieces, while others — like dust conglomerates or some asteroids and planetesimals — are clusters made of loosely bound individual units or particles (Murdoch et al., 2015; Sánchez, 2016; Walsh, 2018; Hestroffer et al., 2019). The larger clusters are held together mainly by gravity, while the small ones are dominated by contact forces.
Symbol     Description Value(s) Units
$R_{i}$ Radius of the $i$-th particle $0.75,1.5,2.25]\times 10^{6}$ cm
$m_{i}$ Mass of the $i$-th particle $0.48,3.8,12.9]\times 10^{19}$ g
$\rho$ Particle density 2.7 g/cm3
@@
$\vec{v}_{i}$ Velocity vector of the $i$-th particle cm/s
$\vec{\omega}_{i}$ Angular velocity vector of the $i$-th particle rad/s
@@
$\vec{F}_{\rm{elastic}}$ Normal elastic force between two particles dyne
$\vec{F}_{i,\rm{friction}}$ The kinetic friction force on particle $i$ dyne
$\vec{F}_{i,\rm{rolling\,friction}}$ The rolling friction force on particle $i$ dyne
$\vec{T}_{ij}$ The torque on particle $i$ due to its interaction with particle $j$ dyne cm
$k$ Spring constant $10^{18}$ dyne/cm
$\mu_{k}$ Coefficient of kinetic friction 0.3
$\mu_{r}$ Coefficient of rolling friction $10^{-3}$
$\delta{t}$ time step of the code 0.04 s
$m_{\rm{p}}$ mass of the projectile (the smaller cluster) g
$m_{\rm{t}}$ mass of the target (the larger cluster) g
$m_{\rm{lf}}$ mass of the largest fragment after collision g
@@
In general, based on observational evidence and previous studies (Wetherill, 1967; Dohnanyi, 1969; Chapman et al., 1978; Farinella et al., 1982; Davis et al., 1985; Michel et al., 2001; Benavidez et al., 2012; Leinhardt & Stewart, 2012; Walsh, 2018), the result of collisions between clusters is not always constructive. Here, by constructive we mean that the largest debris of the collision is more massive than the target cluster [^1] Here and in the following we will call the projectile the cluster with the lower mass, while we will call the target the more massive one. Since our calculations are carried out in the center of mass frame, in which the net momentum is zero, the projectile also has the largest velocity. In case the two rubble piles have the same mass, one of the two will arbitrarily be the projectile and the other the target, if such distinction is needed.. Although higher-velocity collisions are expected to have a higher likelihood of being destructive, the collision outcome is not simply predictable by the energy balance of the collision. Previous work (Farinella et al., 1982; Love & Ahrens, 1996; Asphaug et al., 1998; Durda et al., 2004; Richardson et al., 2009) has indeed shown that even collisions with kinetic energy larger than the binding energy (in absolute value) can be constructive, resulting in a massive bound cluster surrounded by fast debris that carry most of the excess kinetic energy. Additional energy is dissipated in inelastic collisions that increase the thermal energy of the particles. Leinhardt & Stewart (2012), for example, found that even collisions in which the kinetic energy of the center of mass (CoM) is three times higher than the binding energy of the gravitational potential of the clusters can be constructive. The high resilience of granular objects (or clusters) to destructive collisions can be ascribed to several properties that differentiate them from gaseous bodies. Clusters of hard particles obey the dynamics of granular materials. In these structures the pressure gradient is not nearly as isotropic as it is in stars. Instead, the gravity is balanced through contact forces between particles. This results in force chains that propagate collision stresses unevenly among particles. In some configurations, some particles may not be part of any force chains (Clark et al., 2012, 2014; Cheng et al., 2018). Since the forces generated by a collision are propagated through the clusters via contact points, the inelasticity of the particles that compose the cluster allows them to absorb a fraction of the impact energy at every contact point Tanga et al. (2009); Ferrari & Tanga (2020). This causes some of the impact energy to be converted into internal energy, increasing the thermodynamic temperature of the cluster particles. Finally, impacts on porous media create scattered shocks whose energy is confined by the voids, leading to lower levels of disruption (Asphaug et al., 1998).
@@
In this paper we present a new code to study the outcome of subsonic collisions between clusters with masses in the asteroid and planetesimal range. We first present the code (Section 2) and compare its results with the outcome of previous studies (Section 3.1). We then use the capabilities of our code to study the effect of collisions on the particle size distribution of the pre- and post-collision clusters, as well as the effect of rotation of the projectile and target clusters (Section 3). We also analyze the dynamics of the collisions to tease out which particle size is mostly involved in propagating forces during the impact. We finally discuss our results and compare with previous studies (Section 5).
@@ 2 The DECCO Code
In this section we present our soft-sphere, discrete element code DECCO (Discrete Element Cosmic COllision) that we used in the calculations presented in section 3. This type of simulation was originally proposed for granular dynamics by Cundall (1971) (see also Cundall & Strack 1979) and is often used to model granular physics systems (e.g., Hart et al. 1988; Mahmood & Elektorowicz 2016; Sánchez & Scheeres 2011). However, alternative methods exist, such as contact dynamics, quasi-statics, and event driven methods (see, e.g., Radjai & Dubois 2013 for a review). Our code uses the soft-sphere method to compute the binary collision between particle pairs [^2] Collisions between multiple particles are also considered as the cumulative effect of multiple binary collisions., and includes the effects of inelastic collisions as well as kinetic and rotational friction between particles. At the present stage of development, DECCO only includes gravity as the attractive force and is therefore best suited for collisions between macroscopic bodies, such as rubble-pile asteroids or planetesimals. Since DECCO is based on individual particle-particle interactions, it is applicable for impact velocities that are lower than the speed of sound in the particles. Hydrodynamic codes would be needed for faster collisions (see, e.g., Michel et al. 2001; Leinhardt & Stewart 2012; Michel et al. 2020). DECCO also calculates explicitly all binary interactions, yielding increased accuracy with respect to approximate methods (e.g., tree codes and mean field approximation) at the cost of $O(n^{2})$ computational cost. The half-step Verlet method (Verlet, 1967) is used to integrate the equations of motion of the particles. The equations of the method read:
$x[i+1]$ $=$ $x[i]+v[i]dt+\frac{1}{2}a[i]dt^{2}$
$v\left[i+\frac{1}{2}\right]$ $=$ $v[i]+\frac{1}{2}a[i]dt$ (1)
$v[i+1]$ $=$ $v\left[i+\frac{1}{2}\right]+\frac{1}{2}a[i+1]dt$
where $x[i]$ is position at time step $i$, $v[i]$ is velocity at time step $i$, and $a[i]$ is acceleration calculated at position $x[i]$.
@@ 2.1 Free particles
When there is no contact between two particles, i.e., the distance $r_{ij}$ from their centers is larger than the sum of the two particle radii ($r_{ij}>R_{i}+R_{j}$), the only acting force is gravitational attraction. In this case the force that the $i$-th particle feels because of the presence of the $j$-th particle is:
@@
Figure 2: Example outcome of the post-collision sub-cluster identification algorithm. The figure shows the post-collision and re-collapse configuration from an $\eta=3$ collision between two identical rubble piles (5000 particles each). The area shown is where the largest debris are found. Particles belonging to dynamically bound sub-clusters with at least 50 particle members are color coded. Dark gray particles are free and unbound. There are six sub-clusters identified in the figure. The two biggest ones (blue and green) are almost identical due to the symmetry of the collision.
@@ 2.2 Radial contact forces
The implementation of DECCO presented here includes the elastic repulsion between two touching particles as the only radial contact force. To take into account the inelastic behavior of real materials, DECCO utilizes two different values of Hook’s constant: a larger value $k_{\rm{in}}$ for the incoming trajectory (particles approaching each other) and a smaller value $k_{\rm{out}}$ for the outgoing trajectory. This methodology is a simplified version of the linear hysteretic model (see, e.g., Walton & Braun 1986; Luding 1998, 2007; An & Tannant 2007) that ensures a constant coefficient of restitution. Writing conservation of energy for the inbound and outbound motion we have:
$\frac{1}{2}mv_{\rm{in}}^{2}$ $=$ $\frac{1}{2}k_{\rm{in}}\Delta x$
$\frac{1}{2}mv_{\rm{out}}^{2}$ $=$ $\frac{1}{2}k_{\rm{out}}\Delta x$
where $\Delta x$ is the amount of compression of the spring. Taking the ratio of the two equations above one gets:
$\left(\frac{v_{\rm{in}}}{v_{\rm{out}}}\right)^{2}=\frac{k_{\rm{in}}}{k_{\rm{out}}}$
from which a constant coefficient of restitution $\epsilon$ is derived as:
$\epsilon=\frac{v_{\rm{out}}}{v_{\rm{in}}}=\sqrt{\frac{k_{\rm{out}}}{k_{\rm{in}}}}.$ (4)
@@
$\vec{F}_{\rm{elastic}}=k\frac{R_{i}+R_{j}-r_{ij}}{2}\hat{r}_{ij}$ (5)
where $r_{ij}$ is the distance between particles $i$ and $j$, $R_{i}$ and $R_{j}$ are the respective radii of particles $i$ and $j$, and the factor $\frac{R_{i}+R_{j}-r_{ij}}{2}$ measures the deformation of the spring, with the factor $\frac{1}{2}$ taking into account that there are two springs in series  [^3] An alternative method to simulate the particle-particle bouncing force is by the use of a Hertzian spring (see, e.g., Landau & Lifshitz 1959, or Sánchez & Scheeres 2011, 2012, 2014 for examples in the context of asteroids). For the kind of simulations discussed in this paper, we found that a Hertzian spring model causes increased computational cost without an appreciable difference in the outcome.. The elastic constant $k$ is either $k_{\rm{in}}$ or $k_{\rm{out}}$, depending on whether the particles are approaching or receding, respectively. When particles are colliding, this additional surface normal force is added to the gravitational force for the current timestep to simulate the collision ‘bounce’.
@@ 2.3 Restitution
As introduced above, Kinetic energy is removed from particles in collision by a coefficient of restitution directly applied through a change in the elasticity constant, $k$. This method allows for scale-free, constant restitution, which is particularly important since we aim at simulating objects made up of particles of different sizes.
@@ 2.4 Friction
[…]
@@
[…]
where $\mu_{k}$ is the coefficient of friction. We emphasize that DECCO does not include a treatment for static friction nor twisting friction. As such, it does not satisfy Coulomb’s friction criterion and therefore cannot be used to study long timescale stability of the clusters. In addition, the decrease in particle size due to compression was neglected. This can affect the calculation of the surface velocities as well as the torques between particles. Given the imposed limit on spring compression discussed below for code stability, these effects are expected to be minor.
In addition to sliding friction, the code includes rolling friction (e.g., Lunding 2008; Schwartz et al. 2012; Santos et al. 2020), which is calculated as:
[…]
where $\mu_{r}$ is the coefficient of rolling friction. It has to be noted that modeling rolling friction is a complicated task since any deviation from a pure spherical particle shape would affect their rolling behavior much more than their sliding behavior. To understand the reason, let us consider a small particle rolling on a big one. If the small particle has a complex shape (without spherical or cylindrical symmetry), the rolling motion would cause the distance between the two centers of mass to change with time, a kind of motion that requires more energy than the rolling of spherical particles which does not involve any changes in distance (see also Zhou et al. 2017 for further discussion).
The total friction force (sliding plus rolling) affects the center-of-mass linear velocities of the involved particles and contributes a torque that also affects their rotational velocities. The total torque on particle $i$ due to the contact with particle $j$ is then:
@@
and the equation for the evolution of angular velocity of particle $i$ is:
$\frac{2}{5}m_{i}R_{i}^{2}\frac{d\omega_{i}}{dt}=$ (11)
$=$ $\sum_{i\neq j}\frac{R_{i}}{R_{i}+R_{j}}\vec{r}_{ij}\times(\vec{F}_{i,friction}+\vec{F}_{i,rollingFriction})\qquad$
where the inertia of the spherical particles has been assumed to be ${I}=\frac{2}{5}\,m_{i}R_{i}^{2}$. The friction force is equal and opposite for particle $j$, so the necessary information to move both interacting particles is known from a single calculation. The angular acceleration of a particle due to surface friction is found by dividing the torque on a particle by its moment of inertia.